An Investment Club placed a portion of its funds into a 9% annual simple interest account and the remainder into an 8% annual simple interest account. The amount of interest for one year was $860. If the amounts placed in each account had been revered, the interest earned would have been $840. How much was placed in each account?
Answer by Earlsdon(6287)

Let x = the amount placed in the 9% account and y = the amount placed in the 8% account. Change the percentages to their decimal equivalents and write the equations relating the interests earned on each account to the total interest earned.
0.09x+0.08y = $860 Now write the equation for the reverse.
0.08x+0.09y = $840 Multiply each equation by 100 to clear the decimals.
9x+8y = 86000
8x+9y = 84000 Now multiply the top equation by 8 and the bottom one by 9 to get:
72x+64y = 688000
72x+81y = 756000 Subtract the top equation from the bottom one to get:
17y = 68000 Divide by 17.
y = 4000 Now substitute this into one of the previous equations and solve for x.
9x+8y = 86000 Substitute y = 4000.
9x+32000 = 86000 Subtract 32000.
9x = 54000 Finally, divide by 9.
x = 6000
The club placed $6,000.00 into the 9% account and $4,000.00 into the 8% account.